a rectangular garden is surrounding by a walk of uniform width. The area of a garden is 180 sq yards. If the dimensions of the garden plus the walk are 16 yds by 24 yds, what is the width of the walk?
Answer by rwm(914)

16*24=384
x*y=180
(x+w)*(y+w)=384
x+w=16
y+w=24
w=6
x=10
y=18